30 cm3 of 0.1 M Al (NO3)3 SOLUTION IS RECTED WITH 100cm3 of 0.15M of NaOH solution. Which reactant is in excess, and by now much ?

  • A. NaOH solution by 70 cm3
  • B. NaOH solution, by 60 cm3
  • C. NaOH solution by 40 cm3
  • D. Al(NO3)3 solution by 20 cm3
  • E. Al(NO3)3 solution by 10 cm3
1
Expert's answer
2022-02-28T11:03:50-0500

Al(NO3)3 + 3NaOH = 3NaNO3 + Al(OH)3

CM = n / V

n = CM x V

n (Al(NO3)3) = 0.1 x 0.03 = 0.003 mol

n (NaOH) = 0.15 x 0.1 = 0.015 mol

According to the eqation, the required n (NaOH) = 3 x n (Al(NO3)3) = 3 x 0.003 = 0.009 mol

Therefore, n (NaOH) is in excess of 0.015 - 0.009 = 0.006 mol

V (NaOH)excess = 0.006 / 0.15 = 0.04 L = 40 ml


Answer: C. NaOH solution by 40 cm3